The cheetah is the fastest land mammal and is highly specialized to run down prey. The cheetah often exceeds speeds of 60 mil pls explain about C.
Answer #1

Solution:-

Data Mean 59.57429 Standard Error 0.734656 Median 59.2 Mode 57.8 Standard Deviation 4.346286 Sample Variance 18.8902 Kurtosis

Interval for containing observations within one standard deviation is 59.575 + 4.346 or (55.23, 63.92)

Interval for containing observations within two standard deviation is 59.575 + 2*4.346 or (50.88, 68.27)

Interval for containing observations within three standard deviation is 59.575 + 3*4.346 or (46.54, 72.61)

a)

Number of observations within one standard deviations = 27

Total number of observations = 35

Percentage of observation within one standard deviation = (27/35)*100

Percentage of observation within one standard deviation is 77.14%

b)

Number of observations within two standard deviations = 34

Total number of observations = 35

Percentage of observation within two standard deviation = (34/35)*100

Percentage of observation within two standard deviation = 97.14%

c)

Number of observations within three standard deviations 34

Total number of observations = 35

Percentage of observation within three standard deviation = (34/35)*100

Percentage of observation within three standard deviation is 97.14%